OpenStudy (is3535):

A camping goods store has 5 models of waterproof flashlights. The 5 models sell for $23, $29, $14, $26, and $18. What is the mean absolute deviation of the prices? Round to the nearest cent, if necessary.

OpenStudy (is3535):

(23 + 29 + 14 + 26 + 18) / 5 = 110 / 5 = 22

OpenStudy (is3535):

23 - 22 = 1 29 - 22 = 7 22 - 14 = 8 26 - 22 = 4 22 - 18 = 4

OpenStudy (is3535):

(1 + 7 + 8 + 4 + 4) / 5 = 24 / 5 = 4.8

OpenStudy (is3535):

@mathstudent55

OpenStudy (is3535):

@TheOnlyDarkKnight

Vocaloid (vocaloid):

yea, your answer looks fine to me

OpenStudy (is3535):

4.8